OpenStudy (anonymous):

Write the answer to the nearest tenth 449/0.42 so would I be doing this 449/0.42=1069.047619?? then 1069.000000??? HELP please ;-(

OpenStudy (anonymous):

\[\frac{ 449 }{0.42}=\frac{ 449 }{\frac{ 42 }{ 100 } }=\frac{ 449*100 }{42 }=\frac{ 44900 } { 42 }\] now you can solve it.

OpenStudy (anonymous):

im sorry how did you even start that??? im lost sorry

OpenStudy (anonymous):

im looking in my book for better examples I just see a lot of things but thanks for the help..

OpenStudy (anonymous):

change 0.42=42/100

OpenStudy (anonymous):

449/.42 is equal to 1069.047619 the nearest tenth is the first digit after the decimal point. since the second digit after the decimal point is equal to less than 5, you just truncate to the first digit after the decimal place. your answer would be 1069.0 Ok so I was right, I just found this, but again thanks for trying to help ...